An engineer uses a river (minimum temperature of 4°C) to cool his electrical plant, which heats water with a boiler operating at 220°C. His boiler uses 420 kW of energy to heat water to steam. What is the maximum power output under ideal circumstances?

Multiple Choice

  • 416 kW

  • 330 kW

  • 230 kW

  • 184 kW
Answer #1

(Page 1) Tempenature of sink = Te = 4 c = 273 +4 Tc=277k. Temperatuese 04_source - Th= 220°c - In=973+290 Tha 493k Heal eneeg